# Few-Shot Scenario Testing for Autonomous Vehicles Based on Neighborhood Coverage and Similarity

### v1

@@ III-C Optimizing for a Few-Shot Scenario Set
Eventually with the information of SMs set $\mathcal{M}$, the gradient descent method is conducted to search for a optimal test set $\mathcal{X}(s,n)$ and dynamically adjust coverage of samples $x_{i}$ in each iteration. By directly applying the error upper bound in Eq. (8), we write the objective function for optimization as
@@
where $\tilde{\mu}_{i}$ calculated with Eq. (11)-(13).
With the extended upper bound of error in Eq. (10), we can see that $\Delta{\mu}=E_{p}[\Delta P_{i}(A|X)]$ is the unknown overall gap between real AV model and SM set. This item is irrelevant to the estimation strategy $f$ or scenario samples $\mathcal{X}_{s,n}$ and is impossible to be eliminated. Therefore, we focus on the additional part $|\Delta\tilde{\mu}|$ in Eq. (10), which can be written as
$|\Delta\tilde{\mu}|=|\sum_{i=1}^{n}{\Delta P(A|x_{i})w(x_{i};\mathcal{X}_{s,n})}|.$ (15)
@@
This fluctuation estimator is the summation of differences between $x_{i}$ and scenarios in its coverage set weighted by the similarity measurement and exposure frequency. It is practically effective that if a scenario has significant difference with another similar one, the underlying uncertainty should be taken into consideration.
@@
Note that $w_{\mathcal{M}}$ is our confidence parameter on the prior knowledge provided by SMs. If the AV model is predictable with the assistance of SMs, $w_{\mathcal{M}}$ is supposed to be set to +$\infty$+ ∞ and the objective function in Eq. (17) degrade into Eq. (14). The gradient descent method is also applicable with this additional form.
By leveraging scenario coverage and similarity, FST method is capable of selecting optimal test set with an upper bound of error. When the unknown gap between AV and SMs can not be neglected, additional error caused by scenarios selected can also be bounded. Because the formulation targets at the optimized testing error given specific hyper-parameter $n$ and utilizes global information to select test set with strongest generalization ability among AV models, FST is possible to tackle the inaccuracy issue with small testing numbers. As the randomness is restricted in a finite range, it can be used to measure whether the testing error is acceptable within fixed number $n$ of scenarios.
@@ IV-A Cut-in Scenario
where $R$ and $\dot{R}$ denote respectively the range and range rate between background vehicle (BV) and AV at the moment of lane change. Consider the event $A$ as accidents, if the AV fails to speed down to avoid BV in scenario $x$ such that the distance between two vehicles break the threshold $d=s_{BV}-s_{AV}<d_{th}$, the AV would crash with BV. If accidents happen we have $P(A|x)=1$ and otherwise $P(A|x)=0$. By extracting real world driving behaviors from NDD, we construct the exposure frequency $p(x)$ of each scenario variable at cut-in moments. Therefore $P(A)$ denotes the overall accident rate of AV. On the basis of NDD, the scenario space is restricted within
@@
By initializing specific cut-in scenario and simulating actions for AV and BV for a sufficient period of time, $P(A|x)$ and accident rate for vehicle models can be obtained.
@@ IV-B Experiment Settings
=$\mathcal{M}=\{m^{\prime}|m^{\prime}=\sum_{i=1}^{4}{c_{i}m_{i}},\ c_{i}\geq 0,\ \forall i,\ \sum_{i=1}^{4}{c_{i}}=1\}.$ (20)
The performance of SMs is shown in Fig. (2). The separate lines are the accident boundaries of 4 SMs between $P(A|x)=0$ and $P(A|x)=1$. Accidents appear in the left side of boundaries with smaller range and range rate and the overall accident rate of SMs varies from $4.6\times 10^{-4}$ to $4.9\times 10^{-3}$. The exposure frequency of scenarios in NDD is illustrated with the saturability of background color. We can see that most scenarios with large exposure frequency $p(x)$ are accident-free. This results in the rarity of accident events.
Figure 2: Illustration of basic testing methods and SMs. Scenarios with smaller range and range rate in the left side of boundary would encounter accidents in simulation. The saturability of background demonstrate the exposure frequency in NDD.
The AV model for test is set to another group of IDM parameters with the accident rate of $3.0\times 10^{-4}$ and BV keeps a constant velocity after changing lane. As the simulation in our experiment is a complex temporal process, it is hard to fitting the accident performance of AV in all scenarios with 4 SMs and we have $m^{*}\notin\mathcal{M}$. Therefore, we set the confidence parameter $w_{\mathcal{M}}=1$.
@@
As comparisons, we applied CMC testing in NDE and the uniform sampling method random quasi-Monte Carlo (RQMC)[26] to test the AV model. The results of 100 testing scenarios generated by NDE and uniform sampling are also shown in Fig. (2). Because of the rarity of accident events, almost all scenarios generated in NDE concentrate in unchallenging areas, which make the testing result effectless. RQMC method selects scenarios uniformly in the scenarios space with randomness. The scenarios information is extracted evenly to evaluate the performace of AV. With a small number of test, scenarios with high risks or accidents might be tested.
With the purpose of validating FST method with different size of test set and examining the error bound, we set $n=5,10,20$ as the restricted numbers of test. With each testing cost, the baseline methods and FST method were repeated for 100 times.
@@ IV-C Evaluation Results
We use an example of 20 scenarios sampled by FST in Fig. (3) to illustrate the scenario samples selection and coverage division strategy. We use different colors to distinguish the coverage set of samples with the saturability hinting the exposure frequency. The accident boundary of AV model under test is shown by the red line. We can see that the coverage set is large for samples in regions where the scenarios are not challenging enough. It means that only a smaller number of tests are sufficient to verify the performance of AVs. In regions where the SMs exhibit disparate performances, more scenarios are sampled and the coverage of each sample is relatively small. This ensures the generalization ability of FST method among distinct AVs in the prior knowledge set $\mathcal{M}$. Moreover, the accident boundary of SMs and AV matched roughly with the boundary of different coverage set, which allows a smaller error bound and higher accuracy for FST method.
@@
The testing results of NDE, uniform sampling and FST are shown in Fig. (4). For NDE testing, most experiments yield an overall accident rate of 00 while several experiments produce huge testing errors. It indicates that the accurate performance of AV can hardly be obtained with a small test set in NDE. Uniform sampling method achieves a more accurate evaluation result than NDE but in some experiments the error is large. Compared to the other two methods, the maximum errors of FST method using 5, 10, 20 samples are all bounded in a certain scope (the blue dashed line in Fig. (3)) around the real accident rate rate of AV.
@@
The average error and variance of 3 methods are shown in TABLE I. We can see that the FST method surpasses the common baseline methods significantly in all experiments and indices. Furthermore, the accuracy of FST is less impaired than the other two methods when the testing number is small. This remarkable feature shows the efficiency and reliability of FST method to testing with restricted costs or searching for minimum test set.
@@ V Conclusion
In this paper, we propose the few-shot testing method to tackle the problem of quantifying AV performance with a strictly limited number of tests. By utilizing the scenario neighborhood coverage and similarity on prior information of models, we iteratively search for a small scenario set to extract scenarios information with the strongest generalization ability. A theoretical error bound can also be estimated with FST to measure whether the accuracy of testing result is acceptable for a fixed size of test set. Results show that proposed method achieves better accuracy than commonly used baseline methods particularly in case that the limit on testing numbers is strict. In future, more efficient and general methods for the design of scenario weight, coverage and the fluctuation estimator could be developed and our method is possible to be extended to scenarios with higher complexity.

### v2

@@ III-C Optimizing for a Few-Shot Scenario Set
Eventually with the information of SMs set $\mathcal{M}$, the gradient descent method is conducted to search for a optimal test set $\mathcal{X}_{s,n}$ and dynamically adjust coverage of samples $x_{i}$ in each iteration. By directly applying the error upper bound in Eq. (8), we write the objective function for optimization as
@@
where $\tilde{\mu}_{i}$ is calculated with Eq. (11)-(14).
With the extended upper bound of error in Eq. (10), we can see that $\Delta{\mu}=\mathbb{E}_{p}[\Delta P_{i}(A|X)]$ is the unknown overall gap between real AV model and SM set. This item is irrelevant to the estimation strategy $f$ or scenario samples $\mathcal{X}_{s,n}$ and is impossible to be eliminated. Therefore, we focus on the additional part $|\Delta\tilde{\mu}|$ in Eq. (10), which can be written as
$|\Delta\tilde{\mu}|=\left|\sum_{i=1}^{n}{\Delta P(A|x_{i})w(x_{i};\mathcal{X}_{s,n})}\right|.$ (16)
@@
This fluctuation estimator is the summation of differences between $x_{i}$ and scenarios in its coverage set weighted by the similarity measurement and exposure frequency. It is practically effective that if a scenario has significant differences with another one that is identified as similar, the underlying uncertainty should be taken into consideration.
@@
Note that $w_{\mathcal{M}}$ is our confidence parameter on the prior knowledge provided by SMs. If the AV model is predictable with the assistance of SMs, $w_{\mathcal{M}}$ is supposed to be set to +$\infty$+ ∞ and the objective function in Eq. (18) degrades into Eq. (15). The gradient descent method is also applicable with this additional form.
By leveraging scenario coverage and similarity, FST method is capable of selecting optimal test set with an upper bound of error. When the unknown gap between AV and SMs can not be neglected, additional error caused by scenarios selected can also be controlled. Because the formulation targets at the optimized testing error given specific hyper-parameter $n$ and utilizes global information to select test set with strongest generalization ability among AV models, FST is possible to tackle the inaccuracy issue with small testing numbers. As the randomness is restricted in a finite range, it can be used to measure whether the testing error is acceptable within fixed number $n$ of scenarios.
@@ IV-A Cut-in Scenario
where $R$ and $\dot{R}$ denote the range and range rate respectively between the background vehicle (BV) and AV at the moment of lane change. Considering the event $A$ as crashes, if the AV fails to speed down to avoid BV in scenario $x$ such that the distance between two vehicles break the threshold $d=s_{BV}-s_{AV}<d_{th}$, the AV would crash with BV. If crashes happen we have $P(A|x)=1$ and otherwise $P(A|x)=0$. By extracting real world driving behaviors from NDD, we construct the exposure frequency $p(x)$ of each scenario variable at cut-in moments. Therefore $P(A)$ denotes the overall crash rate of AV. On the basis of NDD, the scenario space is restricted within
@@
By initializing specific cut-in scenario and simulating actions for AV and BV for a sufficient period of time, $P(A|x)$ and crash rate for vehicle models can be tested.
@@ IV-B Experiment Settings
=$\mathcal{M}=\left\{m^{\prime}|m^{\prime}=\sum_{i=1}^{4}{c_{i}m_{i}},\ c_{i}\geq 0,\ \forall i,\ \sum_{i=1}^{4}{c_{i}}=1\right\}.$ (21)
The performances of SMs are shown in Fig. 2. The separate lines are the crash boundaries of 4 SMs between $P(A|x)=0$ and $P(A|x)=1$. Accidents appear in the left side of boundaries with smaller range and range rate and the overall crash rates of SMs vary from $4.6\times 10^{-4}$ to $4.9\times 10^{-3}$. The exposure frequency of scenarios in NDD is illustrated with the saturability of background color. We can see that most scenarios with large exposure frequency $p(x)$ are crash-free. This results in the rarity of crash events.
Figure 2: Illustration of basic testing methods and SMs. Scenarios with smaller range and range rate in the left side of boundary would encounter crashes in simulation. The saturability of background demonstrate the exposure frequency in NDD.
The AV model for test is set to another group of IDM parameters with the crash rate of $3.0\times 10^{-4}$ and BV keeps a constant velocity after changing lane. As the simulation in our experiment is a complex temporal process, it is hard to fitting the exact crash performance of AV in all scenarios with 4 SMs and we have $m^{*}\notin\mathcal{M}$. Therefore, we set the confidence parameter $w_{\mathcal{M}}=1$.
@@
As comparisons, we applied CMC testing in NDE and the uniform sampling method random quasi-Monte Carlo (RQMC)[31] to test the AV model. The results of 100 testing scenarios generated by NDE and uniform sampling are also shown in Fig. 2. Because of the rarity of crash events, almost all scenarios generated in NDE concentrate in unchallenging areas, which make the testing and evaluation result effectless. RQMC selects scenarios uniformly in the scenarios space with randomness. The scenarios information is extracted evenly to evaluate the performace of AV. With a small number of test, scenarios with high risks might be tested.
With the purpose of verifying FST method with diverse sizes of test set and examining the error bound, we set $n=5,10,20$ as the restricted numbers of test. With each number of testing scenarios, the baseline methods and FST method were repeated for 100 times.
@@ IV-C Evaluation Results
We use an example of 20 scenarios sampled by FST in Fig. 3 to illustrate the scenario samples selection and coverage division strategy. The different colors is used to distinguish the coverage set of samples with the saturability hinting the exposure frequency. The crash boundary of AV model under test is shown with the red line. We can see that the coverage set is large for samples in regions where the scenarios are not challenging enough. It means that only a smaller number of tests are sufficient to verify the performance of AVs. In regions where the SMs exhibit disparate performances, more scenarios are sampled and the coverage of each sample is relatively small. This ensures the generalization ability of FST method among distinct AVs in the prior knowledge set $\mathcal{M}$. Moreover, the crash boundary of SMs and AV matches roughly with the boundary of different coverage set, which allows a smaller error bound and higher accuracy for FST method.
@@
The testing and evaluation results of NDE, uniform sampling and FST are shown in Fig. 4. For NDE testing, most experiments yield an overall crash rate of 00 while several experiments produce huge testing errors. It indicates that the accurate performance of AV can hardly be obtained with a small test set in NDE. Uniform sampling method achieves a more accurate evaluation result than NDE but in some experiments the error is large. Compared to the other two methods, the maximum errors of FST method using 5, 10, 20 samples are all bounded in a certain scope (the blue dashed line in Fig. 4 around the real crash rate of AV.
@@
The average error and variance of 3 methods are shown in TABLE I. The FST method surpasses the common baseline methods significantly in all experiments and indices. Furthermore, the accuracy of FST is less impaired than the other two methods when the number of testing scenarios is smaller. This remarkable feature shows the efficiency and reliability of FST method to testing with restricted budgets or searching for minimum test set.
@@ V Conclusion
In this paper, we propose the few-shot testing method to tackle the problem of quantifying AV performance with a strictly limited number of tests. By utilizing the scenario neighborhood coverage and similarity on prior information of models, we iteratively search for a small scenario set to extract scenarios information with the strongest generalization ability. A theoretical error bound can also be estimated with FST to measure whether the accuracy of testing and evaluation result is acceptable for a fixed size of test set. Results show that the proposed method achieves better accuracy than commonly used baseline methods particularly in case that the limit on testing numbers is stricter. In future, more efficient and general designment of scenario weight, coverage and the fluctuation estimator could be developed. Besides, the application of FST method on more complex scenarios is also a potential direction of further studies.